OpenStudy (anonymous):

what number is rational btween 0 and 0.5

OpenStudy (anonymous):

There are infinitely many rational numbers between any two points on the number line.

OpenStudy (anonymous):

but i need to know exactly a number

OpenStudy (anonymous):

well what is the rational form of .5?

OpenStudy (anonymous):

idk?

OpenStudy (anonymous):

\[.5 = \frac{.5}{1} = \frac{5}{10} = \frac{5*1}{5*2} = \frac{1}{2}\]

OpenStudy (anonymous):

oo so the answer would be 1/2